ROTATIONAL SPECTRUM OF N_{2}–CO: THE ORTHO-N_{2} AND PARA-N_{2}

\underline{Y. KAWASHIMA}, Department of Applied Chemistry, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, JAPAN ; Y. OHSHIMA, Department of Chemistry, Kyoto University, Kyoto 606- 8502, JAPAN ; Y. ENDO, Graduate School of Art and Science, University of Tokyo 153-8902, JAPAN.

Observations of the infrared spectra of N_{2}–CO by diode laser spectroscopy have been reported previously^{a} ^{b}. The present study is an observation of the rotational spectra of ^{14}N_{2}–CO, ^{15}N^{14}N–CO, and ^{15}N_{2}–CO using a Fourier transform microwave spectrometer. Two sets for the J = 2-1, 3-2, and 4-3 transitions are observed for ^{14}N_{2}–CO and ^{15}N_{2}–CO, while only one set for ^{15}N^{14}N–CO. The hyperfine structures are observed for the complexes containing ^{14}N nucleus. The first set is assigned to the K=0 state of ortho-N_{2} and the second to the K=1 state of para-N_{2},based on the infrared analysis and eQq splittings. The transitions with the K=1 state are stronger than those with the K=0 state for ^{14}N_{2}–CO complex, while the reversed is observed for ^{15}N_{2}–CO. In the case of ^{15}N^{14}N–CO only lowest transition is observed because of lack of symmetry in the complex. The nuclear electric quadrupole coupling constant of the ^{14}N nucleus is much smaller than related complexes. The value of the van der Waals stretching frequency for N_{2}–CO is calculated to be 19 cm^{−1} from the rotational and the centrifugal distortion constants obtained. We concluded that both moieties, N_{2} and CO, are rotating almost freely in the complex N_{2}–CO.

aY. Kawashima and K. Nishizawa, Chem. Phys. Lett. 249, 87 (1996)

bY. Xu and A. R. W. McKellar, J. Chem. Phys. 104, 2488 (1996).